a^2+b^2=52 and a^2-b^2= 20 Solve for b
Answer by venugopalramana(3286)

a^2+b^2=52 .......................I
and a^2-b^2= 20............................II
Solve for b
EQN.I-EQN.II GIVES
A^2+B^2-(A^2-B^2)=52-20=32
A^2+B^2-A^2+B^2=32
2B^2=32
B^2=32/2=16
B=+4 OR -4